# Complexity of Linear Subsequences of $k$-Automatic Sequences

### v2

@@ 5 Construction by Büchi Arithmetic
Given the $m$-state msd-first DFAO generating $(h(i))_{i\geq 0}$, the msd-first DFAO generating $(h(i+j))_{i,j\geq 0}$ is computed in $O(m^{2}\log m)$ time.
@@
Next, we project the $z$ component away and use subset construction to construct the DFAO $M$. On input $x\times y$, the automaton $M$ reaches a state consisting of $0,p_{0}]$ and $1,p_{1}]$ where $p_{0},p_{1}\in Q_{h}$ and we have $M(x,y)=\tau_{h}(p_{0})$.
@@
We continue by constructing automaton $M_{n}$ recognizing one of the most basic relations, that is, $n[x]_{k}=[z]_{k}$ where $n$ is a fixed constant. We have $M_{n}(x,z)=1$ if and only if $n[x]_{k}=[z]_{k}$. An obvious implementation of this (and the one used by Walnut) uses the so-called “binary method” , also called “Russian multiplication”, which is implemented recursively.
The base case is $n=1$. Otherwise, either $n$ is even or $n$ is odd.
If $n$ is even, we recursively compute a DFA $M_{n/2}(x,y)$ recognizing $y]_{k}=(n/2)[x]_{k}$ and then use the first-order expression
$\exists y\ M_{n/2}(x,y)\,\wedge\,M_{\text{add}}(y,y,z),$
$\exists y\ M_{n-1}(x,y)\,\wedge\,M_{\text{add}}(x,y,z),$
which is similarly translated into an automaton.
In both cases, the translation could conceivably generate a nondeterministic automaton (by the projection). In this case, the subset construction is used to determinize the resulting automaton. Thus, at least in principle, this particular translation mechanism could take exponential time. However, we show that in this case, the subset construction does not cause an exponential blowup in the number of states. For now, let us ignore the cost of the minimization steps.
Theorem 23.
Let $n\geq 1$ be a fixed constant. The msd-first DFA $M_{n}$ for recognizing $y]_{k}=n[x]_{k}$ constructed by the binary method has $n$ states corresponding to differences $y]_{k}-n[x]_{k}$ in the range [0,n1][0,n-1].
Proof.
We prove the result by induction on $n$.
Base case: $n=1$.
Case 1: $n$ is even.
By the induction hypothesis, we have a DFA $M_{n/2}$ with $n/2$ states corresponding to differences $y]_{k}-(n/2)[x]_{k}$ in the range [0,n/21][0,n/2-1]. In this case, we use the formula
$\exists y\ M_{n/2}(x,y)\,\wedge\,M_{\text{add}}(y,y,z).$
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n-1]\},$
$\delta(d,[a,c])$ $=kd+c-na,$
$q_{0}$ $=0,$
$A$ $=\{q_{0}\}.$
Case 2: $n$ is odd.
$\exists y\ M_{n-1}(x,y)\,\wedge\,M_{\text{add}}(x,y,z).$
Next, we project away the $y$ component. The resulting automaton is nondeterministic, and we use subset construction to create a deterministic automaton $M_{n}$. Consider a state from $M_{n}$ reachable on $x\times z$. This state consists of all pairs $[q,p]$ from $M^{\prime}_{n}$ such that $z]_{k}-n[x]_{k}=q+p$. As mentioned above, if $z]_{k}-n[x]_{k}$ is 0 or n1n-1, then there is only one such pair; otherwise, there are two such pairs. Furthermore, no $[q,p]$ appears in more than one state of $M_{n}$. So $M_{n}$ has $n$ states and the determinization did not cause a blow up in the number of states. The DFA $M_{n}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n-1]\},$
$\delta(d,[a,e])$ $=kd+e-na,$
$q_{0}$ $=0,$
$A$ $=\{q_{0}\}.$
Theorem 24.
Let $n\geq 1$ be a fixed constant. An msd-first automaton for recognizing $y]_{k}=n[x]_{k}$ on input $x\times y$ can be computed in $O(n\log n)$ time.
Proof.
We use $T(n)$ to show the time required to recursively create $M_{n}$. We now analyze $T(n)$ based on the information from the proof of Theorem 23.
In Case 2, the number of transitions traversed to obtain $M_{n}$ is as follows: $4$ transitions for $M_{\text{add}}$, $2(n-1)k^{3}$ transitions for the product construction of $M_{n-1}$ and $M_{\text{add}}$, $2(n-1)k^{3}$ transitions for projecting away the $y$ component, $nk^{2}$ transitions for the determinized automaton, giving a total of $O(n)$ transitions in addition to the number of transitions traversed to create $M_{n-1}$.
$\begin{cases}T(n)=O(n\log n)+T(n/2),&\text{if $n\bmod 2=0,$}\\ T(n)=O(n\log n)+T(n-1),&\text{if $n\bmod 2=1$}.\end{cases}$
So we have $T(n)=O(n\log n)$. ∎
Theorem 25.
Let $c\geq 0$ be a fixed constant. The msd-first automaton recognizing the relation $x]_{k}+c=[z]_{k}$ can be computed in $O(\log c)$ time.
Proof.
Consider the $O(\log c)$ state automaton $M_{=c}=(Q_{=c},\Sigma_{k},\delta_{=c},q_{0,=c},A_{=c})$ recognizing $y]_{k}=c$ on input $y$. This automaton ignores leading zeros and compares input to $(c)_{k}$ letter by letter. Each state indicates how many letters from $y$ have been matched with $(c)_{k}$ except for leading zeros. We have $M_{=c}(y)=1$ if and only if $y]_{k}=c$. The only inputs $y$ that do not lead to the dead state are of the form $0^{*}y^{\prime}$ where $y^{\prime}$ is a (potentially length-0) prefix of $(c)_{k}$.
We can use the following expression to create automaton $M_{c}$ recognizing $x]_{k}+c=[z]_{k}$ on input $x\times z$.
$\exists y\ M_{\text{add}}(x,y,z)\,\wedge\,M_{=c}(y).$
First, we apply the product construction to $M_{\text{add}}$ and $M_{=c}$ such that the $M_{\text{add}}$ part processes $x\times y\times z$ and the $M_{=c}$ part processes $y$. The resulting automaton $M^{\prime}_{c}$ has $O(\log c)$ states. Let $\operatorname{Pre}_{k}(c,i)$ denote the first $i$ letters in $(c)_{k}$. On input $x\times y\times z$ where $y$ is $0^{*}y^{\prime}$ and $y^{\prime}$ is a prefix of $(c)_{k}$ the automaton $M^{\prime}_{c}$ leads to a state $[q,p]$ where $q\in\{0,1\}$ is from $M_{\text{add}}$ and $p=|y^{\prime}|$ is from $M_{=c}$ indicating how many letters from $y$ have been matched with $(c)_{k}$. Therefore, we have $q=[z]_{k}-([x]_{k}+[y]_{k})$ and $\operatorname{Pre}_{k}(c,p)]_{k}=[y]_{k}$. So $z]_{k}-[x]_{k}=q+[\operatorname{Pre}_{k}(c,p)]_{k}$ and to each state $[q,p]$ in $M^{\prime}_{c}$ a difference $z]_{k}-[x]_{k}$ can be attributed.
Then, we remove the $y$ component from transitions and determinize the automaton to obtain $M_{c}$. Consider a (non-dead) state reachable on input $x\times z$ in $M_{c}$. This state consists of all $[q,p]$ such that $z]_{k}-[x]_{k}=q+[\operatorname{Pre}_{k}(c,p)]_{k}$. Since $q\in\{0,1\}$, there are at most two such $[q,p]$. Furthermore, each $[q,p]$ from $M^{\prime}_{c}$ appears in at most one state in $M_{c}$. Therefore, there are $O(\log c)$ states in $M_{c}$. The resulting automaton $M_{c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,c+1]\},$
$\delta(d^{\prime},[a,e])$ $=kd^{\prime}+e-a,$
$q_{0}$ $=0,$
$A$ $=\{c\}.$
Note that not all $d\in[0,c+1]$ are reachable.
So it takes $O(\log c)$ time to create automaton $M_{c}$ recognizing $x]_{k}+c=[z]_{k}$ on input $x\times z$. ∎
Theorem 26.
Let $n\geq 1$, $c\geq 0$ be fixed constants. The msd-first automaton recognizing $n[x]_{k}+c=[z]_{k}$ on input $x\times z$ can be computed in time $O((n+c)\log(c)\log(nc))$.
Proof.
Now we want to create an automaton $M_{n,c}$ recognizing $n[x]_{k}+c=[z]_{k}$.
Let $M_{n}$ be the automaton from Theorem 24 and let $M_{c}$ be the automaton from Theorem 25. We have $M_{n}(x,y)=1$ if and only if $n[x]_{k}=[y]_{k}$ and $M_{c}(y,z)=1$ if and only if $y]_{k}+c=[z]_{k}$. So we can use the following expression to create $M_{n,c}$:
$\exists y\ M_{n}(x,y)\,\wedge\,M_{c}(y,z).$
Consider an input $x\times y\times z$ for $M^{\prime}_{n,c}$ leading to some state $[q,p]$ where $q$ is from $M_{n}$ and $p$ is from $M_{c}$. By the construction of $M^{\prime}_{n,c}$, we know that $q=[y]_{k}-n[x]_{k}$ and $p=[z]_{k}-[y]_{k}$. So we have $q+p=[z]_{k}-n[x]_{k}$. Therefore, to each state in automaton $M^{\prime}_{n,c}$ a different $z]_{k}-n[x]_{k}$ can be attributed. Furthermore, since $q\in\{0,\ldots,n-1\}$ and $p\in[0,c+1]$, the attributed differences are in the range $[0,n+c]$. If we choose a difference $d\in[0,n+c]$, then we can exactly tell what combinations of $[q,p]$ correspond to $d$ (if any) and no $[q,p]$ corresponds to more than one $d$.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n+c]\},$
$\delta(d,[a,e])$ $=kd+e-na,$
$q_{0}$ $=0,$
$A$ $=\{c\}.$
Taking into account the minimization steps and the time spent on creating $M_{n}$ and $M_{c}$ based on the proof of theorems 24 and 25, the time spent on creating $M_{n,c}$ is $O(n\log(c)\log(nc)+(n+c)\log(n+c))$. So for large enough values of $n$ and $c$, the time spent on creating $M_{n,c}$ is $O((n+c)\log(c)\log(nc))$. ∎
Theorem 27.
Let $n\geq 1$, $c\geq 0$ be fixed constants. Given an $m$-state msd-first DFAO generating $(h(i))_{i\geq 0}$, we can compute the msd-first DFAO for $(h(ni+c))_{i\geq 0}$ in $O(m^{2}(n+c)(\log(m)+\log(c)\log(nc))$ time.
Proof.
We first use the product construction for $M_{n,c}$ and $M_{h}$ such that the $M_{n,c}$ part processes the input $x\times y$ and the $M_{h}$ part processes the input $y$ to get automaton $M^{\prime}$. Consider a state $[q,p]$ reachable on $x\times y$ in $M^{\prime}$ where $q$ is from $M_{n,c}$ and $p$ is from $M_{h}$. We have $q=[y]_{k}-n[x]_{k}\in[0,n+c]$ and $p=\delta_{h}(q_{0,h},y)$. This automaton has $m(n+c+1)$ states.
Next, we project the $y$ component away and use subset construction to construct the DFAO $M$. On input $x$ the automaton reaches a state consisting of all pairs $[q,p]$ where $q\in[0,n+c]$ and $p\in Q_{h}$. Among these pairs there is one pair $q^{\prime},p^{\prime}]$ such that $q^{\prime}=c$ and we have $M(x)=\tau_{h}(p^{\prime})$. The DFAO $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ can be formally defined as follows. Let $i=n+c$.
$Q$ $=\{[[0,p_{0}],\ldots,[i,p_{i}]]:p_{j}\in Q_{h}\},$
$\delta([[0,p^{\prime}_{0}],\ldots,[i,p^{\prime}_{i}]],a)$ $=[[0,p_{0}],\ldots,[i,p_{i}]]$
$q_{0}$ $=[[0,q_{0,h}],\ldots,[i,\delta_{h}(q_{0,h},i)]],$
$\Delta$ $=\Delta_{h},$
$\tau([[0,p_{0}],\ldots,[i,p_{i}]])$ $=\{\tau_{h}(p_{j})\ :\ j=c\}.$
We need to know that the transition function is well-defined and has the correct output. This can be done similar to the proof Theorem 10.
In each state of $M$, the first components of the pairs are consecutive differences in the range $[0,n+c]$ and the $p$ values are a subsequence of the interior sequence $(h^{\prime}(i))_{i\geq 0}$. Therefore, there are $O(m^{2}(n+c))$ states in $M$. Taking into account the minimization steps and the time spent on creating $M_{n,c}$ based on the proof of Theorem 26, creating $M$ given $M_{h}$ takes $O(n\log(c)\log(nc)+m^{2}(n+c)\log(m)+m^{2}(n+c)\log(n+c)$ time. So for large enough values of $n$ and $c$, the time spent on creating $M$ is $O(m^{2}(n+c)(\log(m)+\log(c)\log(nc))$. ∎

### v3

@@ 5 Construction by Büchi Arithmetic
Let $c\geq 0$ be a fixed constant. The automaton recognizing the relation $x]_{k}=c$ with msd-first input can be computed in $O((\log^{2}c)(\log\log c))$ time.
Proof.
We recursively create the automaton $M_{=c}$.
The base case is $c=0$ and the automaton has $1$ state.
If $c$ is even, we recursively compute a minimal DFA $M_{=(c/2)}$ recognizing $y]_{k}=c/2$ on input $y$ and then use the first-order expression
$\exists y\ M_{=(c/2)}(y)\land M_{\text{add}}(y,y,x)$
which is translated into an automaton. We know the automaton for $M_{\text{add}}$ has $2$ states and by Lemma 23, the automaton $M_{=(c/2)}$ has $\lfloor\log_{k}(c/2)\rfloor$ + 2 states. First we apply the product construction to $M_{=(c/2)}$ and $M_{\text{add}}$ such that the $M_{=(c/2)}$ part processes $y$ and the $M_{\text{add}}$ part processes $y\times y\times x$; the resulting automaton has $2(\lfloor\log_{k}(c/2)\rfloor+2)$ states. Each state in the product automaton is of the form $[q,p]$ where on input $x\times y$, $q$ is a state from the minimal automaton $M_{=(c/2)}$ representing the number of letters from $(c/2)_{k}$ matched with $y$ by Lemma 23 and $p\in\{0,1\}$ is a state from $M_{\text{add}}$ equal to $x]_{k}-2[y]_{k}$. Then we apply the $\exists y$ quantifier by removing the component corresponding to $y$ from the transitions and we get an NFA. If on input $x$ this NFA reaches $[q,p]$, from $q$ and $p$ we can determine how many letters from $x$ match $(c)_{k}$. So each state in this NFA corresponds to a state in the automaton in the proof of Lemma 23 for constant $c$. Then we do a subset construction to create a DFA from the NFA and the resulting DFA is the automaton from Lemma 23 with $O(\log c)$ states.
If $c$ is odd, we recursively compute a DFA $M_{=(c-1)}$ recognizing $y]_{k}=c-1$ on input $y$ and then use the first-order expression
$\exists y\ M_{(c-1)}(y)\land M_{\text{add}}(y,1,x)$
which is translated into an automaton. Following a construction similar to the one described for the case where $c$ is even, we get the DFA from Lemma 23 for constant $c$ with $O(\log c)$ states.
Taking minimization steps into account, we get the following recursive formula for the time complexity:
$T(c)=\begin{cases}O((\log c)(\log\log c))+T(c/2),&\text{if $c\bmod 2=0$};\\ O((\log c)(\log\log c))+T(c-1),&\text{if $c\bmod 2=1$}.\end{cases}$
So the total time is $O((\log^{2}c)(\log\log c))$. ∎
Theorem 25.
Let $c\geq 0$ be a fixed constant. The automaton recognizing the relation $x]_{k}+c=[z]_{k}$ with msd-first input can be computed in $O((\log^{2}c)(\log\log c))$ time.
Proof.
Consider the $O(\log c)$ state automaton $M_{=c}$ from Theorem 24 recognizing $y]_{k}=c$ on input $y$. We have $M_{=c}(y)=1$ if and only if $y]_{k}=c$.
We can use the following expression to create the automaton $M_{c}$ recognizing $x]_{k}+c=[z]_{k}$ on input $x\times z$.
$\exists y\ M_{\text{add}}(x,y,z)\,\wedge\,M_{=c}(y).$
First, we apply the $\land$ by product construction for $M_{\text{add}}$ and $M_{=c}$ such that the $M_{\text{add}}$ part processes $x\times y\times z$ and the $M_{=c}$ part processes $y$. The resulting automaton $M^{\prime}_{c}$ has $O(\log c)$ states. Let $\operatorname{Pref}_{k}(c,i)$ denote the first $i$ letters in $(c)_{k}$. On input $x\times y\times z$ where $y$ is $0^{*}y^{\prime}$ and $y^{\prime}$ is a prefix of $(c)_{k}$, the automaton $M^{\prime}_{c}$ leads to a state $[q,p]$ where $q\in\{0,1\}$ is from $M_{\text{add}}$ and $p=|y^{\prime}|$ is from $M_{=c}$ indicating how many letters from $y$ have been matched with $(c)_{k}$. Therefore, we have $q=[z]_{k}-([x]_{k}+[y]_{k})$ and $\operatorname{Pref}_{k}(c,p)]_{k}=[y]_{k}$. So $z]_{k}-[x]_{k}=q+[\operatorname{Pref}_{k}(c,p)]_{k}$ and to each state $[q,p]$ in $M^{\prime}_{c}$ a difference $z]_{k}-[x]_{k}$ can be attributed.
Then to implement the $\exists y$ quantifier, we remove the $y$ component from all transitions. We determinize the automaton to obtain $M_{c}$. Consider a (non-dead) state reachable on input $x\times z$ in $M_{c}$. This state consists of all $[q,p]$ such that $z]_{k}-[x]_{k}=q+[\operatorname{Pref}_{k}(c,p)]_{k}$. Since $q\in\{0,1\}$, there are at most two such $[q,p]$. Furthermore, each $[q,p]$ from $M^{\prime}_{c}$ appears in at most one state in $M_{c}$. Therefore, there are $O(\log c)$ states in $M_{c}$. The resulting automaton $M_{c}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,c+1]\},$
$\delta(d^{\prime},[a,e])$ $=kd^{\prime}+e-a,$
$q_{0}$ $=0,$
$A$ $=\{c\}.$
Note that not all $d\in[0,c+1]$ are reachable from the initial state.
Considering the time spent on creating $M_{=c}$ and the minimization steps, it takes $O((\log^{2}c)(\log\log c))$ time to create the automaton $M_{c}$ recognizing $x]_{k}+c=[z]_{k}$ on input $x\times z$. ∎
We continue by constructing the automaton $M_{n}$ recognizing $n[x]_{k}=[z]_{k}$ on input $x\times z$ where $n$ is a fixed constant. We have $M_{n}(x,z)=1$ if and only if $n[x]_{k}=[z]_{k}$. There are several possible recursions for creating this automaton; for example, we could use the following:
  • $\exists y\ M_{n-1}(x,y)\,\wedge\,M_{\text{add}}(x,y,z)$.
  • $\exists y_{1},y_{2}\ M_{\lfloor n/2\rfloor}(x,y_{1})\,\wedge\,M_{\lceil n/2\rceil}(x,y_{2})\,\wedge\,M_{\text{add}}(y_{1},y_{2},z)$.
However, there is a third recursive implementation that is more efficient than the two above, and we describe it now. The base case is $n=1$. Otherwise, either $n$ is even or $n$ is odd.
If $n$ is even, we recursively compute a DFA $M_{n/2}$ recognizing $y]_{k}=(n/2)[x]_{k}$ on input $x\times y$ and then use the first-order expression
$\exists y\ M_{n/2}(x,y)\,\wedge\,M_{\text{add}}(y,y,z),$
$\exists y\ M_{n-1}(x,y)\,\wedge\,M_{\text{add}}(x,y,z),$
which is similarly translated into an automaton.
In both cases, the translation could conceivably generate a nondeterministic automaton (by the projection). In this case, the subset construction is used to determinize the resulting automaton. Thus, at least in principle, this particular translation mechanism could take exponential time. However, we show that in this case, the subset construction does not cause an exponential blowup in the number of states.
Theorem 26.
Let $n\geq 1$ be a fixed constant. The DFA $M_{n}$ with msd-first input for recognizing $y]_{k}=n[x]_{k}$ constructed recursively has $n$ states corresponding to differences $y]_{k}-n[x]_{k}$ in the range [0,n1][0,n-1].
Proof.
We prove the result by induction on $n$.
Base case: $n=1$.
Case 1: $n$ is even.
By the induction hypothesis, we have a DFA $M_{n/2}$ with $n/2$ states corresponding to differences $y]_{k}-(n/2)[x]_{k}$ in the range [0,(n/2)1][0,(n/2)-1]. In this case, we use the formula
$\exists y\ M_{n/2}(x,y)\,\wedge\,M_{\text{add}}(y,y,z).$
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n-1]\},$
$\delta(d,[a,e])$ $=kd+e-na,$
$q_{0}$ $=0,$
$A$ $=\{q_{0}\}.$
Case 2: $n$ is odd.
$\exists y\ M_{n-1}(x,y)\,\wedge\,M_{\text{add}}(x,y,z).$
Next, we project away the $y$ component. The resulting automaton is nondeterministic, and we use the subset construction to create a deterministic automaton $M_{n}$. Consider a state from $M_{n}$ reachable on $x\times z$. This state consists of all pairs $[q,p]$ from $M^{\prime}_{n}$ such that $z]_{k}-n[x]_{k}=q+p$. As mentioned above, if $z]_{k}-n[x]_{k}$ is 0 or n1n-1, then there is only one such pair; otherwise, there are two such pairs. Furthermore, no $[q,p]$ appears in more than one state of $M_{n}$. So $M_{n}$ has $n$ states and the determinization did not cause a blow up in the number of states. The DFA $M_{n}=(Q,\Sigma^{2}_{k},\delta,q_{0},A)$ can be formally defined as follows.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n-1]\},$
$\delta(d,[a,e])$ $=kd+e-na,$
$q_{0}$ $=0,$
$A$ $=\{q_{0}\}.$
Theorem 27.
Let $n\geq 1$ be a fixed constant. The automaton for recognizing $y]_{k}=n[x]_{k}$ on msd-first input can be computed in $O(n\log^{2}n)$ time.
Proof.
Let $T(n)$ be the time used by the algorithm above to create $M_{n}$. We now analyze $T(n)$ based on the information from the proof of Theorem 26.
In Case 2, the number of transitions traversed to obtain $M_{n}$ is as follows: $4$ transitions for $M_{\text{add}}$, $2(n-1)k^{3}$ transitions for the product construction of $M_{n-1}$ and $M_{\text{add}}$, $2(n-1)k^{3}$ transitions for projecting away the $y$ component, $nk^{2}$ transitions for the determinized automaton, giving a total of $O(n)$ transitions in addition to the number of transitions to create $M_{n-1}$.
$T(n)=\begin{cases}O(n\log n)+T(n/2),&\text{if $n\bmod 2=0$};\\ O(n\log n)+T(n-1),&\text{if $n\bmod 2=1$}.\end{cases}$
So we have $T(n)=O(n\log^{2}n)$. ∎
Theorem 28.
Let $n\geq 1$, $c\geq 0$ be fixed constants. For large enough values of $n$ and $c$, the automaton recognizing the relation $n[x]_{k}+c=[z]_{k}$ with msd-first input can be computed in $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c)))$ time.
Proof.
Now we want to create an automaton $M_{n,c}$ recognizing $n[x]_{k}+c=[z]_{k}$.
Let $M_{n}$ be the automaton from Theorem 27 for constant $n$ and let $M_{c}$ be the automaton from Theorem 25 for constant $c$. We have $M_{n}(x,y)=1$ if and only if $n[x]_{k}=[y]_{k}$ and $M_{c}(y,z)=1$ if and only if $y]_{k}+c=[z]_{k}$. So we can use the following expression to create $M_{n,c}$:
$\exists y\ M_{n}(x,y)\,\wedge\,M_{c}(y,z).$
Consider an input $x\times y\times z$ for $M^{\prime}_{n,c}$ leading to some state $[q,p]$ where $q$ is from $M_{n}$ and $p$ is from $M_{c}$. By the construction of $M^{\prime}_{n,c}$, we know that $q=[y]_{k}-n[x]_{k}$ and $p=[z]_{k}-[y]_{k}$. So we have $q+p=[z]_{k}-n[x]_{k}$. Therefore, to each state in the automaton $M^{\prime}_{n,c}$ a different $z]_{k}-n[x]_{k}$ can be attributed. Furthermore, since $q\in\{0,\ldots,n-1\}$ and $p\in[0,c+1]$, the attributed differences are in the range $[0,n+c]$. If we choose a difference $d\in[0,n+c]$, then we can exactly tell what combinations of $[q,p]$ correspond to $d$ (if any) and no $[q,p]$ corresponds to more than one $d$.
$Q$ $=\{d\in\mathbb{N}\ :\ d\in[0,n+c]\},$
$\delta(d,[a,e])$ $=kd+e-na,$
$q_{0}$ $=0,$
$A$ $=\{c\}.$
Taking into account the minimization steps and the time spent on creating $M_{n}$ and $M_{c}$ based on the proof of Theorems 27 and 25, the time spent on creating $M_{n,c}$ is $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c))+(n+c)\log(n+c))$. So for large enough values of $n$ and $c$, the time spent on creating $M_{n,c}$ is $M_{n,c}$ is $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c)))$. ∎
Theorem 29.
@@
Let $n\geq 1$, $c\geq 0$ be fixed constants. For msd-first input, given an $m$-state DFAO generating $(h(i))_{i\geq 0}$, for large enough values of $n$ and $c$, we can compute the DFAO generating $(h(ni+c))_{i\geq 0}$ in $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c))+m^{2}(n+c)\log(m^{2}(n+c)))$ time.
Proof.
We first use the product construction for $M_{n,c}$ and $M_{h}$ such that the $M_{n,c}$ part processes the input $x\times y$ and the $M_{h}$ part processes the input $y$ to get the automaton $M^{\prime}$. Consider a state $[q,p]$ reachable on $x\times y$ in $M^{\prime}$ where $q$ is from $M_{n,c}$ and $p$ is from $M_{h}$. We have $q=[y]_{k}-n[x]_{k}\in[0,n+c]$ and $p=\delta_{h}(q_{0,h},y)$. This automaton has $m(n+c+1)$ states.
Next, we project the $y$ component away and use the subset construction to construct the DFAO $M$. On input $x$ the automaton reaches a state consisting of all pairs $[q,p]$ where $q\in[0,n+c]$ and $p=\delta_{h}(q_{0,h},(n[x]_{k}+q)_{k})$. Among these pairs there is one pair $q^{\prime},p^{\prime}]$ such that $q^{\prime}=c$ and we have $M(x)=\tau_{h}(p^{\prime})$. The DFAO $M=(Q,\Sigma_{k},\delta,q_{0},\Delta,\tau)$ can be formally defined as follows. Let $l=n+c$.
$Q$ $=\{[[0,p_{0}],\ldots,[l,p_{l}]]:p_{j}\in Q_{h}\},$
$\delta([[0,p^{\prime}_{0}],\ldots,[l,p^{\prime}_{l}]],a)$ $=[[0,p_{0}],\ldots,[c,p_{l}]]$
$q_{0}$ $=[[0,q_{0,h}],\ldots,[l,\delta_{h}(q_{0,h},(l)_{k})]],$
$\Delta$ $=\Delta_{h},$
$\tau([[0,p_{0}],\ldots,[l,p_{l}]])$ $=\{\tau_{h}(p_{j})\ :\ j=c\}.$
We need to know that the transition function is well-defined and has the correct output. The proof is similar to the proof of Theorem 10.
In each state of $M$, the first components of the pairs are consecutive differences in the range $[0,n+c]$ and the $p$ values are a subsequence of the interior sequence $(h^{\prime}(i))_{i\geq 0}$. Therefore, there are $O(m^{2}(n+c))$ states in $M$. Taking into account the minimization steps and the time spent on creating $M_{n,c}$ based on the proof of Theorem 28, for large enough values of $n$ and $c$, creating $M$ given $M_{h}$ takes $O(\log^{2}(c)\log\log(c)+n\log^{2}(n)+n\log(c)\log(n\log(c))+m^{2}(n+c)\log(m^{2}(n+c)))$ time. ∎
Theorem 30.
For msd-first input, given the $m$-state DFAO generating $(h(i))_{i\geq 0}$, the DFAO generating $(h(i+j))_{i,j\geq 0}$ is computed in $O(m^{2}\log m)$ time.
@@
Next, we project the $z$ component away and use the subset construction to construct the DFAO $M$. On input $x\times y$, the automaton $M$ reaches a state consisting of $0,p_{0}]$ and $1,p_{1}]$ where $p_{q}=\delta_{h}(q_{0,h},(q+[x]_{k}+[y]_{k})_{k})$ and we have $M(x,y)=\tau_{h}(p_{0})$.